OpenStudy (anonymous):

Nick works two jobs to pay for college. He tutors for $15 per hour and also works as a bag boy for $8 per hour. Due to his class and study schedule, Nick is only able to work up to 20 hours per week, but must earn at least $150 per week. If t represents the number of hours Nick tutors and b represents the number of hours he works as a bag boy, which system of inequalities represents this scenario? t + b >= 20 15t + 8b = 150 t + b <= 20 15t + 8b >= 150 t + b <= 20 15t + 8b <= 150 None of the systems shown represent this scenario.

OpenStudy (anonymous):

@ZeHanz can u help with this one too ?

OpenStudy (anonymous):

i have this and two others left

OpenStudy (zehanz):

OK, I will give it a try! t is the number of tutor hours, b is the number of bag hours. The total number of hours worked is t+b, can be no more than 20 His earnings with these activities will be 15t+8b, must at least be 150. Could you translate this to the right inequalities now?

OpenStudy (anonymous):

not into the exact inequalities but i think i would be dealt with t + b <= 20 15t + 8b >= 150 either that or none of the systems shown

OpenStudy (zehanz):

Looks OK to me!

OpenStudy (anonymous):

so which one then ? @ZeHanz

OpenStudy (zehanz):

You chose option B yourself, so I didn't think you needed more help. So, It's B, it's B, it's B! (you may sing it)

OpenStudy (anonymous):

lol thanks i was sure if it was b or d

OpenStudy (anonymous):

The graph below is represented by which system of inequalities? -x + 3y < 9 y > -2x + 1 -x + 3y <= 9 y > -2x + 1 -x + 3y < 9 y -2x + 1 None of these systems represent the graph shown.

OpenStudy (zehanz):

Have a little confidence in your reasoning. You are doing fine!

OpenStudy (anonymous):